OpenStudy (theviper):

U can use this formula to find the value of 'x' in Quadratic Equations It's so useful

OpenStudy (theviper):

\[\Huge{\color{red}{\frac{-b \pm \sqrt{b^2-4ac}}{2a}}}\]

OpenStudy (theviper):

use this formula for equations like \[ax^2\pm bx \pm c\]

OpenStudy (anonymous):

cause you the viper http://www.youtube.com/watch?v=TKFH_zh4gY0

OpenStudy (cwrw238):

is there a question with this?

OpenStudy (theviper):

e.g, 5x^2-3x-1 here a=5, b=-3 & c=-1

OpenStudy (cwrw238):

right

OpenStudy (theviper):

Putting these values in formula, we will get two values.

OpenStudy (theviper):

for x*

OpenStudy (cwrw238):

yes - in this case you will get 2 values for x

OpenStudy (theviper):

first, \[x=\frac{-5+\sqrt{29}}{10}\] second, \[x=\frac{-5-\sqrt{29}}{10}\]

OpenStudy (cwrw238):

no its -b in the numerator - that is -(-3) = 3

OpenStudy (cwrw238):

not -5

OpenStudy (cwrw238):

ty lol

OpenStudy (theviper):

yes u got that I was trying if anyone could catch this before giving me a medal haha & u deserved a medal:)

OpenStudy (theviper):

so\[x=\frac{3+\sqrt{29}}{10}\]or\[x=\frac{3-\sqrt{29}}{10}\]

OpenStudy (cwrw238):

yea

OpenStudy (theviper):

haha thanx sir @satellite73 :)

OpenStudy (theviper):

@satellite73 sir plz plz plz change ur pic to this:- http://upload.wikimedia.org/wikipedia/commons/8/8d/GPS_Satellite_NASA_art-iif.jpg

OpenStudy (theviper):

it will match ur name @satellite73

OpenStudy (anonymous):

@TheViper, I'm glad you wanted to share your joy with the quadratic formula. :)

OpenStudy (theviper):

thanx @Limitless

OpenStudy (anonymous):

the quadratic formula is very useful :)

OpenStudy (theviper):

yes right

OpenStudy (anonymous):

Anyone up for the Cubic Formula (for a leading coefficient of 1)? \[Ax^3+Bx^2+Cx+D=0\] Let \(a=\frac{B}{A}\), \(b=\frac{C}{A}\), and \(c=\frac{D}{A}.\) Then, \[ \begin{eqnarray*} r_1 & = & -\frac{a}{3} + \left(\frac{-2a^3 + 9ab - 27c + \sqrt{(2a^3-9ab+27c)^2 + 4(-a^2+3b)^3}}{54}\right)^{1/3} \\ & & {} + \left(\frac{-2a^3 + 9ab - 27c - \sqrt{(2a^3-9ab+27c)^2 + 4(-a^2+3b)^3}}{54}\right)^{1/3} \\ r_2 & = & -\frac{a}{3} - \frac{1+i\sqrt{3}}{2} \left(\frac{-2a^3 + 9ab - 27c + \sqrt{(2a^3-9ab+27c)^2 + 4(-a^2+3b)^3}}{54}\right)^{1/3} \\ & & {} + \frac{-1+i\sqrt{3}}{2} \left(\frac{-2a^3 + 9ab - 27c - \sqrt{(2a^3-9ab+27c)^2 + 4(-a^2+3b)^3}}{54}\right)^{1/3} \\ r_3 & = & -\frac{a}{3} + \frac{-1+i\sqrt{3}}{2} \left(\frac{-2a^3 + 9ab - 27c + \sqrt{(2a^3-9ab+27c)^2 + 4(-a^2+3b)^3}}{54}\right)^{1/3} \\ & & {} - \frac{1+i\sqrt{3}}{2} \left(\frac{-2a^3 + 9ab - 27c - \sqrt{(2a^3-9ab+27c)^2 + 4(-a^2+3b)^3}}{54}\right)^{1/3}. \end{eqnarray*} \]

OpenStudy (theviper):

wow

OpenStudy (theviper):

thanx

OpenStudy (anonymous):

source: http://planetmath.org/?method=src&id=1407&op=getobj&from=objects

OpenStudy (theviper):

no limit to give answers

OpenStudy (lgbasallote):

lol @Limitless =))))